OpenStudy (christos):

Find the amplitude and the phase shift of the function f(x) = - 3sin(2x + π/2)

OpenStudy (anonymous):

using the general equation \[f(x)=asin(bx-c)+d a is the amplitude, it measures how high and low the function goes aka max and mins. In otherwords if you had amplitude 3 , it would mean that the graph would max at 3 and min a -3

OpenStudy (anonymous):

phase shift is when x-c =0

OpenStudy (anonymous):

so for amplitude you have a=-3 so amplitude equates to -3. This means that the sine function is a mirror image of the regular function. It starts downwards rather than upwards. as for phase shift you need to be careful with b. If it's 1 , then you can simply say x-c=0 or x=c. another example x-(-c)=0 x+c=0 x=-c

OpenStudy (anonymous):

when you have a b that is not 1 i simply pull it out so now you get \[-3sin(2(x+\frac{\pi}{4}))\]

OpenStudy (anonymous):

so, now you have \[x+\frac{\pi}{4}=0\] \[x=\frac{-\pi}{4}\]

OpenStudy (anonymous):

that is your phase shift

OpenStudy (christos):

ty!!!